Let f : R → R be defined by f(x) = (x3 + 1)/2

a. Prove that f is bijective

b. Determine f -1 (x) and f o f o f -1


1
Expert's answer
2022-05-25T17:08:47-0400

a. Let "f(x_1)=f(x_2)." It means that


"\dfrac{x_1^3+1}{2}=\dfrac{x_2^3+1}{2}""x_1^3=x_2^3""(x_1-x_2)(x_1^2+x_1x_2+x_3^2)=0""x_1-x_2=0""x_1=x_2"

The function "f(x)=\dfrac{x^3+1}{2}" is bijective (one-to-one ) from "\R" to "\R."


b.


"f(x)=\dfrac{x^3+1}{2}, x\in \R""y=\dfrac{x^3+1}{2}"

Change "x" and "y"

"x=\dfrac{y^3+1}{2}"

Solve for "y"

"y^3=2x-1"




"y=\sqrt[3]{2x-1}"

Then



"f^{-1}(x)=\sqrt[3]{2x-1}"


"f\circ f^{-1}=\dfrac{(\sqrt[3]{2x-1})^3+1}{2}=x"


"f\circ f\circ f^{-1}=\dfrac{x^3+1}{2}"